OpenStudy (artemisxrpg):

Two pulleys, one with radius 2 in. and the other with radius 8 in., are connected by a belt. If the 2 in. pulley is caused to rotate at 3 revolutions per min. determine the revolutions per min of the 8 in. pulley.

OpenStudy (anonymous):

I think this would be the equation.. c = 2 pi r a = 2 pi 2 a = 12.566371 b = 2 pi 8 b = 50.265482 revb/reva = (a/b) revb = reva(a/b)

OpenStudy (anonymous):

2*pi*2*3 = 2*pi*8*x........... Solve for x.

OpenStudy (artemisxrpg):

Can you explain a little better?

OpenStudy (anonymous):

The rope goes the same speed wherever it is. On the first pulley that is 2 pi (2) (3/minute) = 2 pi (8) (x/minute) for the second pulley x = (2)(3)/8 = 6/8 = 3/4

OpenStudy (artemisxrpg):

Why do you multiply the 1st and 2nd pulley by 2pi?

OpenStudy (artemisxrpg):

or is that suppose to be the circumferences times the rev/min?